OpenStudy (anonymous):

Help :0 -> Mary purchased a 2-entree plate at panda express. The cost of the meal is $4.95. If the tax rate is 8%, what was the amount of her bill?

OpenStudy (saifoo.khan):

2*(4.95)= Cost for 2 meals.

OpenStudy (anonymous):

What?

OpenStudy (saifoo.khan):

Cost for 2 meals = 2*(4.95) = $9.9 for 2 meals. Now 8% tax on it, so, = 9.9 * 0.08 = 0.792 tax Final amount = 9.9 +0.792 = Answer